# An Advanced Deep Learning Framework for Ischemic and Hemorrhagic Brain Stroke Diagnosis Using Computed Tomography (CT) Images

### v1

@@ Abstract
Brain stroke is one of the leading causes of mortality and long-term disability worldwide, highlighting the need for precise and fast prediction techniques. Computed Tomography (CT) scan is considered one of the most effective methods for diagnosing brain strokes. The majority of stroke classification techniques rely on a single slice-level prediction mechanism, allowing the radiologist to manually choose the most critical CT slice from the original CT volume. Although clinical evaluations are often used in traditional diagnostic procedures, machine learning (ML) has opened up new avenues for improving stroke diagnosis. To supplement traditional diagnostic techniques, this study investigates the use of machine learning models, specifically concerning the prediction of brain stroke at an early stage utilizing CT scan images. In this research, we proposed a novel approach to brain stroke detection leveraging machine learning techniques, focusing on optimizing classification performance with pre-trained deep learning models and advanced optimization strategies. Pre-trained models, including DenseNet201, InceptionV3, MobileNetV2, ResNet50, and Xception, are utilized for feature extraction. Additionally, we employed feature engineering techniques, including BFO, PCA, and LDA, to enhance models’ performance further. These features are subsequently classified using machine learning algorithms such as SVC, RF, XGB, DT, LR, KNN, and GNB. Our experiments demonstrate that the combination of MobileNetV2, LDA, and SVC achieved the highest classification accuracy of 97.93%, significantly outperforming other model-optimizer-classifier combinations. The results underline the effectiveness of integrating lightweight pre-trained models with robust optimization and classification techniques for brain stroke diagnosis.
@@ 1 Introduction
Brain Stroke is a medical condition that is one of the leading causes of disability and death worldwide and is brought on by a disruption of blood supply to the brain [1]. A brain stroke occurs when a blockage or ruptured blood vessel stops blood flow to a part of the brain, causing oxygen deprivation that damages brain cells. Most of the brain strokes are either ischemic or hemorrhagic [2]. An ischemic stroke occurs when a blood clot or blockage prevents blood flow to the brain, depriving brain tissue of oxygen and nutrients and ultimately resulting in the death of brain cells [3]. A hemorrhagic stroke happens when a weak blood artery ruptures, resulting in internal or external brain hemorrhage. It is frequently associated with high blood pressure or aneurysms. This kind of stroke directly damages the brain due to the pressure of accumulated blood [4].
@@
Brain stroke can be identified using various conventional methods, which mostly rely on clinical information and patient histories, magnetic resonance imaging (MRI), and computed tomography (CT) scans [10]. In recent years, Machine learning (ML) has emerged as an effective instrument in the diagnosis, treatment, and prognosis of diseases, including stroke. Machine learning can identify brain stroke early using both image data and categorical data [11]. However, many researchers in the past utilized Machine Learning to detect brain stroke by using CT scans, MRI, ECG, or categorical data; for instance, [12] used CT scans and deep learning techniques to build an automated and efficient way for classifying brain stroke. [13] developed an ensemble learning model to improve early brain stroke identification using MRI images. They combined MobileNet V3 with LightGBM and CatBoost to extract features. Then, they employed the Random Forest classifier to detect brain stroke. [14] made use of the Kaggle Stroke Prediction categorical dataset and traditional machine learning classifiers to detect brain stroke. [15] used two different datasets containing medical records and MRI images to assess the performance of machine learning, deep learning as well as hybrid models to identify brain stroke. [16] proposed the CNN architectures such as DenseNet121, ResNet50, and VGG16 with MRI image data for better addressing the ongoing challenges of brain stroke diagnosis. [17] utilized two open-source cerebral stroke classification datasets with 48381 combined data samples with six boosting algorithms to classify brain stroke. [18] used raw ECG sensor data to introduce a novel deep-learning model to detect stroke at an early stage, bypassing the necessity of analyzing the frequency domain.
According to the study above, researchers used different kinds of datasets, such as MRI, CT scan, ECG, or Categorical data, to diagnose brain stroke. However, Computed Tomography (CT) Scans and Magnetic Resonance Imaging (MRI) are the most widely used and reliable methods [19]. Deep learning and machine learning hold promise for technological advancement in difficult medical scenarios, especially in brain stroke diagnosis. There are some improvements needed in brain stroke classification using machine learning with CT scan data. This research presents a model that uses a CT scan image dataset to detect brain stroke early and prevent mortality and disability by combining many machine-learning approaches. Our major contribution is as follows.
@@
    This research introduced a novel approach to diagnosing brain stroke using computed tomography images.
@@
    By merging and curating datasets, this research introduced a more robust dataset for brain stroke diagnosis. This dataset addresses potential gaps in publicly available CT scan datasets.
@@
    Several feature extraction methods, including ResNet50, DenseNet201, InceptionV3, Xception and MobileNetV2 were utilized
@@
The remaining portions of our investigation are delineated by the following headings: In Section II, a thorough evaluation of the literature is conducted, and the drawbacks of the existing approaches are examined. The dataset and techniques utilized for machine learning-based Brain Stroke detection are reviewed, and a detailed explanation of data preparation and feature extraction is provided in Section III. Section IV describes the experimental setup and implementation. Section V then provides a critical evaluation of the outcomes of merging several machine learning techniques. The conclusion and future application are outlined in Section VI, which also emphasizes the project’s significance in transforming prenatal care.
@@ 2 Literature Review
Numerous researchers have introduced brain stroke detection models using traditional and machine learning techniques to predict brain strokes. However, this field remains relatively underdeveloped, and ongoing research continues to address its gaps. To gain a clearer understanding of the current state of knowledge in this area, several existing studies that used computed tomography (CT) scans to diagnose brain stroke have been reviewed and presented here for further analysis. A. Abumihsan et al. [20] used a hybrid feature extraction method in conjunction with a convolutional block attention module (CBAM) to present a novel method for ischemic stroke identification from computed tomography (CT) images. The hybrid feature extraction method provides a thorough representation of brain CT images by utilizing the advantages of two pre-trained models, DenseNet121 and MobileNetV3, through feature fusion. A special first-hand dataset (Dataset 1) gathered from the Specialized Private Hospital in Palestine was used to design and assess the suggested methodology. The approach was further tested on a publicly available dataset (Dataset 2) to further show its generalizability and robustness. The findings demonstrate that the suggested model performed exceptionally well, achieving an accuracy of 99.21% on Dataset 1 and 98.73% on Dataset 2. However, this study is limited to only two feature extraction methods, which may restrict feature diversity and model performance. M.M Hossain et al. [21] suggested a novel hybrid ViT-LSTM model to accurately detect stroke using CT scan images. Additionally, they applied Explainable AI (XAI) methods like SHAP, attention mapping, and LIME to provide trustworthy and reasonable predictions and assure clinical relevance. They evaluated the model using BrSCTHD-2023 dataset and the Kaggle brain stroke dataset. The proposed ViT-LSTM model performed the best with 94.55% accuracy with layer normalization and 92.22% accuracy without layer normalization using BrSCTHD-2023 dataset. Besides, it achieved 96.61% accuracy on the Kaggle dataset. Despite this, the hybrid ViT-LSTM model requires higher computational cost, making real-time deployment in clinical settings more challenging. I. Tahyudin et al. [22] developed a deep learning model incorporating Convolutional Block Attention Module (CBAM) with ResNet18 to classify stroke in brain CT images. They used a publicly available dataset that contains 2501 patients brain CT images, where the number of neurological stroke instances were 950 and healthy non-stroke instances were 1551. The CBAM-ResNet18 model outperformed the standard ResNet18 model with 95% validation accuracy. It also obtained an AUC score of 0.99, demonstrating its capability to classify the stroke and non-stroke samples well. Regardless of its outstanding performance, it has some drawbacks, such as generalization issues and limited data size. SK. UmaMaheswaran et al. [23] presented a novel hybrid technique to detect acute strokes from CT scan images. They used hybrid preprocessing approaches to enhance the quality of images. After that, they used discrete wavelet transform, Gabor Filter, and local binary pattern to extract the features from the segmented images. Then, they utilized the dingo optimization algorithm to select the optimal features from the extracted images. Afterward, the selected optimal features were provided to the novel machine learning model as input to detect if the stroke was normal or acute. The proposed XGBoost model achieved an accuracy of 97%, sensitivity of 94%, and 0.015 false positive rate. Then, the proposed XGBoost model’s performance was compared with other machine learning models such as NB, SVM, and Artificial Neural Network. The comparison suggests that the proposed XGBoost model outperformed other models. M.A. Saleem et al. [24] developed detection system using CT brain images combining a genetic algorithm to select the most relevant features and a bidirectional long short-term memory (BiLSTM) model to detect brain strokes at the early stage. The BiLSTM model achieved an accuracy of 96.5%. They also used cross-validation to evaluate the model’s overall performance using precision, recall, F1 score, ROC, and AUC. In addition to that, they compared the model’s performance with classical ML models such as SVM, LR, NB, DT, and RF, where the BiLSTM model outperformed all of them. However, the model’s decision-making process is lacking in explainability. M. Sabir et al. [25] introduced a novel deep convolutional neural network (DCNN) model for early-stage brain stroke detection using CT scan images. The model consists of a feature extractor, a feature fusion module, and a stroke detection module. The proposed model performed better than state-of-the-art models like VGG16, ResNet50, and InceptionV3, achieving higher accuracy, sensitivity, and specificity. In their observation, the proposed model has obtained an accuracy of 96.5% to detect stroke within 6 hours of onset. Despite this, the model was trained with limited data, which might have affected its generalizability. B. Borsos et al. [26] presented a multimodal deep learning model trained on Dutch acute stroke audit data and CT perfusion imaging for predicting the acute ischemic stroke of patients after three months of discharge. They developed TabNet and ResNet-10 models distinctly for each data type, where both the models achieved AUC 0.71 and 0.70, respectively. Later, they built a multimodal DAFT model that outperformed the existing baseline models and similar works by achieving an AUC of 0.75 and an F1 score of 0.80 in spite of using small data records of 98%. Nevertheless, small amounts of data may impact generalizability. Data must be extensive and multi-centered. S. Prasher et al. [27] developed a deep learning model based on EfficientNet-B0 CNN architecture to predict brain strokes from 2515 CT scan images. The proposed model performed impressively, achieving an accuracy of 98.72%. Thus, they demonstrated its ability to help doctors detect brain strokes from CT images. The researchers think that the model might better identify strokes by being trained with more CT images to provide better diagnoses to stroke patients. However, there is a lack of diversity in the data samples.
@@
Table 1: Summary Table for Related Works on Brain Stroke Diagnosis Using CT Scan Images
@@
[26] DAFT
80.00%
Small amounts of data were utilized.
@@
[29] 3D CNN 92.50%
Make use of a small dataset.
[30] UNet 99.57%
Trained with noisy dataset, and class imbalance exists in the data.
@@
C.D. Kulathilake et al. [28] presents a multiclass brain stroke detection model using the NCCT dataset, leveraging two deep learning models based on the DenseNet201 transfer learning architecture. The proposed approach follows a two-step process to identify ischemic stroke and its subtypes. To enhance convergence speed and improve training efficiency, the 1-cycle policy and FastAI framework were integrated into the models. Model 01 achieved a training accuracy of 97.03%, a validation accuracy of 94.50%. Similarly, Model 02 attained a training accuracy of 98.02%, a validation accuracy of 95.67%. The DenseNet201 architecture, combined with FastAI techniques, demonstrated superior performance compared to traditional and task-specific algorithms. However, this research experiment used only one architecture and did not explore other architectures. Y. Ahmed et al. [29] used CT scan images to detect brain strokes by applying a 3D CNN model. They employed data augmentation as well as image preprocessing techniques to improve the quality of the images. The 3D CNN model performed well to detect the strokes in the images, achieving an accuracy of 92.5%. The authors provided the importance of identifying the stroke at an early stage to prevent it. Nonetheless, this study used a small dataset, which can be solved in the future by using a larger dataset to improve the model’s performance. F. Yousaf et al. [30] developed a convolutional neural network (CNN) based model to detect and classify the images of brain tumors and ischemic stroke simultaneously. For this, they created a dataset by merging two public datasets known as BRATS 2015 and ISLES 2015. The CNN model used feature maps to apply fusion to improve the UNET architecture between the encoder blocks to keep the low-level information and address the features that overlap. They split the data into 80:20 ratio. The model showed efficient multi-disease identification and segmentation by achieving 99.56% accuracy, 99.99% specificity, 99.59% precision, and a 99.57% F1-score. Nevertheless, the model was trained with a small and noisy dataset, and class imbalance exists in the data. R. Raj et al. [31] proposed a hybrid framework, namely StrokeViT with AutoML, by combining CNN, Vision Transformers (ViT), and AutoML to improve the diagnosis of stroke from the CT scan images. In this framework, the ViTs were used to extract long-range dependencies, whereas the CNNs were used to capture the local features to improve the slice-wise predictions. It’s noted that the AutoML is applied to generate the predictions at the patient level by integrating slice-level data, where it achieved an accuracy of 87% for the slice-level data as well as successfully classified the patient-level data with 92% accuracy, which surpassed the standalone architectures by an accuracy of 9%. Authors claimed that this technique can adapt for different diseases diagnosed in musculoskeletal and cancer conditions. Nonetheless, the clinical validation is limited in the proposed model. A. Diker et al. [32] introduced a deep learning architecture for categorizing strokes using the Brain Stroke CT scan Dataset. For testing and training purposes, 2501 computed tomography (CT) images of brain strokes were employed in the experimental investigation. This study aimed to categorize brain stroke CT scans as normal and as stroke using a number of well-known pre-trained convolutional neural networks (CNNs), including GoogleNet, AlexNet, VGG-16, VGG-19, and Residual CNN. With an accuracy of 97.06%, VGG-19 obtained the highest score. Though the results are promising, they are limited to a specific and small dataset.
@@ 3.3 Feature Extraction
In this research, we utilize pre-trained deep learning models—DenseNet201, InceptionV3, MobileNetV2, ResNet50, and Xception—to extract discriminative features from CT scan images. The process begins with image preprocessing, where images are resized to match the input dimensions required by each model (e.g., 224×224 pixels for DenseNet201, MobileNetV2, and ResNet50, and 299×299 pixels for InceptionV3 and Xception). The pixel values are then normalized between 0 and 1 to ensure numerical stability and faster convergence during model inference.
@@
3.3.1 Dense Convolutional Network - DensNet201

### v2

# An Efficient Deep Learning Framework for Brain Stroke Diagnosis Using Computed Tomography (CT) Images

@@ Abstract
Brain stroke is a leading cause of mortality and long-term disability worldwide, underscoring the need for precise and rapid prediction techniques. Computed Tomography (CT) scan is considered one of the most effective methods for diagnosing brain strokes. Most stroke classification techniques use a single slice-level prediction mechanism, requiring radiologists to manually select the most critical CT slice from the original CT volume. Although clinical evaluations are often used in traditional diagnostic procedures, machine learning (ML) has opened up new avenues for improving stroke diagnosis. To supplement traditional diagnostic techniques, this study investigates machine learning models for early brain stroke prediction using CT scan images. This research proposes a novel machine learning approach to brain stroke detection, focusing on optimizing classification performance with pre-trained deep learning models and advanced optimization strategies. Pre-trained models, including DenseNet201, InceptionV3, MobileNetV2, ResNet50, and Xception, are used for feature extraction. Feature engineering techniques, including BFO, PCA, and LDA, further enhance model performance. These features are then classified using machine learning algorithms, including SVC, RF, XGB, DT, LR, KNN, and GNB. Our experiments demonstrate that the combination of MobileNetV2, LDA, and SVC achieved the highest classification accuracy of 97.93%, significantly outperforming other model-optimizer-classifier combinations. The results underline the effectiveness of integrating lightweight pre-trained models with robust optimization and classification techniques for brain stroke diagnosis.
@@ 1 Introduction
Brain Stroke is a medical condition, one of the leading causes of disability and death worldwide, and is brought on by a disruption of blood supply to the brain [1]. Brain stroke occurs when a blockage or ruptured blood vessel interrupts blood flow to the brain, causing oxygen deprivation that damages brain cells. Most of the brain strokes are either ischemic or hemorrhagic [2]. An ischemic stroke occurs when a blood clot or blockage prevents blood flow to the brain, depriving brain tissue of oxygen and nutrients and ultimately resulting in the death of brain cells [3]. A hemorrhagic stroke occurs when a weak blood vessel ruptures, causing an internal or external brain hemorrhage. It is frequently associated with high blood pressure or aneurysms. This kind of stroke directly damages the brain due to the pressure of accumulated blood [4].
@@
Brain stroke can be identified using various conventional methods, which mostly rely on clinical information and patient histories, magnetic resonance imaging (MRI), and computed tomography (CT) scans [10]. In recent years, Machine learning (ML) has emerged as an effective instrument in the diagnosis, treatment, and prognosis of diseases, including stroke. Machine learning can identify brain stroke early using both image data and categorical data [11]. However, many researchers in the past utilized Machine Learning to detect brain stroke by using CT scans, MRI, ECG, or categorical data; for instance, [12] [13] used CT images and deep learning techniques for classifying brain stroke. [14] [15] developed deep learning models to improve early brain stroke identification using MRI images. [16] [17] made use of the Stroke Prediction categorical dataset to detect brain stroke. [18] used raw ECG sensor data to detect stroke at an early stage.
According to the study above, researchers used different kinds of datasets, such as MRI, CT scan, ECG, or Categorical data, to diagnose brain stroke. However, Computed Tomography (CT) Scans and Magnetic Resonance Imaging (MRI) are the most widely used and reliable methods [19]. Deep learning and machine learning hold promise for technological advancement in challenging medical scenarios, especially in brain stroke diagnosis. There are some improvements needed in brain stroke classification using machine learning with CT scan data. This research presents a method that uses a CT scan image dataset to detect brain stroke early and prevent mortality and disability by combining many machine-learning approaches. Our major contributions are as follows.
@@
    This research created a more robust dataset for brain stroke diagnosis by merging and curating existing datasets. This dataset addresses potential gaps in publicly available CT scan datasets.
@@
    We utilized several feature extraction methods, including ResNet50, DenseNet201, InceptionV3, Xception, and MobileNetV2.
@@
The remaining sections of this paper are organized as follows: Section II provides a thorough literature review and examines the drawbacks of existing approaches. Section III reviews the dataset and techniques used for machine learning-based brain stroke detection and provides a detailed explanation of data preparation and feature extraction. Section IV describes the experimental setup and implementation. Section V critically evaluates the results of combining several machine learning techniques. Section VI concludes the paper, outlines potential future applications, and emphasizes the project’s significance.
@@ 2 Literature Review
Numerous researchers have introduced brain stroke detection models using traditional and machine learning techniques. However, this field remains relatively underdeveloped, and ongoing research aims to address its limitations. To gain a clearer understanding of the current state of knowledge in this area, several existing studies that used computed tomography (CT) scans to diagnose brain stroke have been reviewed and presented here for further analysis. A. Abumihsan et al. [20] used a hybrid feature extraction method with a convolutional block attention module (CBAM) to introduce a novel method for ischemic stroke identification from CT scan images. The hybrid feature extraction method provides a thorough representation of brain CT images by leveraging the advantages of two pre-trained models, DenseNet121 and MobileNetV3, through feature fusion. A specialized, first-hand dataset (Dataset 1) collected from the Specialized Private Hospital in Palestine was used to design and assess the proposed methodology. The approach was further tested on a publicly available dataset (Dataset 2) to demonstrate its generalizability and robustness. The findings demonstrate that the suggested model performed exceptionally well, achieving an accuracy of 99.21% on Dataset 1 and 98.73% on Dataset 2. However, this study is limited to only two feature extraction methods, which may restrict feature diversity and model performance. M.M Hossain et al. [21] suggested a novel hybrid ViT-LSTM model to accurately detect stroke using CT scan images. Additionally, they applied Explainable AI (XAI) methods like SHAP, attention mapping, and LIME to provide trustworthy and reasonable predictions and assure clinical relevance. They evaluated the model using BrSCTHD-2023 dataset and the Kaggle brain stroke dataset. The proposed ViT-LSTM model performed the best with 94.55% accuracy with layer normalization and 92.22% accuracy without layer normalization using BrSCTHD-2023 dataset. Besides, it achieved 96.61% accuracy on the Kaggle dataset. Despite this, the hybrid ViT-LSTM model requires higher computational cost, making real-time deployment in clinical settings more challenging. I. Tahyudin et al. [22] developed a deep learning model incorporating Convolutional Block Attention Module (CBAM) with ResNet18 to classify stroke in brain CT images. They used a publicly available dataset containing 2501 brain CT images from patients, with 950 neurological stroke instances and 1551 healthy, non-stroke instances. The CBAM-ResNet18 model outperformed the standard ResNet18 model with 95% validation accuracy. It also obtained an AUC score of 0.99, demonstrating its capability to classify the stroke and non-stroke samples well. Regardless of its outstanding performance, it has some drawbacks, such as generalization issues and limited data size.
@@
Table 1: Overview of State-of-the-Art Approaches and Their Limitations on Brain Stroke Diagnosis Using CT Images
@@
SK. UmaMaheswaran et al. [23] presented a novel hybrid technique to detect acute strokes from CT scan images. They used hybrid preprocessing approaches to enhance the quality of images. After that, they used discrete wavelet transform, Gabor Filter, and local binary pattern to extract the features from the segmented images. They then used the dingo optimization algorithm to select optimal features from the extracted images. Afterward, the selected optimal features were input into a novel ML model to detect normal or acute strokes. The proposed XGBoost model achieved an accuracy of 97%. M.A. Saleem et al. [24] developed detection system using CT brain images combining a genetic algorithm to select the most relevant features and a bidirectional long short-term memory (BiLSTM) model to detect brain strokes at the early stage. The BiLSTM model achieved an accuracy of 96.5%. They also used cross-validation to evaluate the model’s overall performance using precision, recall, F1 score, ROC, and AUC. In addition to that, they compared the model’s performance with classical ML models such as SVM, LR, NB, DT, and RF, where the BiLSTM model outperformed all of them. However, the model’s decision-making process is lacking in explainability. M. Sabir et al. [25] introduced a novel deep convolutional neural network (DCNN) model for early-stage brain stroke detection using CT scan images. The model consists of a feature extractor, a feature fusion module, and a stroke detection module. The proposed model performed better than state-of-the-art models like VGG16, ResNet50, and InceptionV3, achieving higher accuracy, sensitivity, and specificity. In their observation, the proposed model has obtained an accuracy of 96.5% to detect stroke within 6 hours of onset. Despite this, the model was trained with limited data, which might have affected its generalizability. S. Prasher et al. [26] developed a deep learning model based on EfficientNet-B0 CNN architecture to predict brain strokes from 2515 CT scan images. The proposed model performed impressively, achieving an accuracy of 98.72%. Thus, they demonstrated its ability to help doctors detect brain strokes from CT images. The researchers think that the model might better identify strokes by being trained with more CT images to provide better diagnoses to stroke patients. However, there is a lack of diversity in the data samples. C.D. Kulathilake et al. [27] presents a multiclass brain stroke detection model using the NCCT dataset, leveraging two deep learning models based on the DenseNet201 transfer learning architecture. The proposed approach follows a two-step process to identify ischemic stroke and its subtypes. To enhance convergence speed and improve training efficiency, the 1-cycle policy and FastAI framework were integrated into the models. Model 01 achieved a training accuracy of 97.03%, a validation accuracy of 94.50%. Similarly, Model 02 attained a training accuracy of 98.02%, a validation accuracy of 95.67%. The DenseNet201 architecture, combined with FastAI techniques, demonstrated superior performance compared to traditional and task-specific algorithms. However, this research experiment used only one architecture and did not explore other architectures. R. Raj et al. [28] proposed a hybrid framework, StrokeViT with AutoML, combining CNNs, Vision Transformers (ViT), and AutoML to improve stroke diagnosis from CT scan images. In this framework, the ViTs were used to extract long-range dependencies, whereas the CNNs were used to capture the local features to improve the slice-wise predictions. AutoML generated patient-level predictions by integrating slice-level data, achieving an accuracy of 87% for the slice-level data as well as successfully classified the patient-level data with 92% accuracy, which surpassed the standalone architectures by an accuracy of 9%. Authors claimed that this technique can adapt for different diseases diagnosed in musculoskeletal and cancer conditions. Nonetheless, the clinical validation is limited in the proposed model. A. Diker et al. [29] introduced a deep learning architecture for categorizing strokes using the Brain Stroke CT scan Dataset. For testing and training purposes, 2501 computed tomography (CT) images of brain strokes were employed in the experimental investigation. This study aimed to categorize brain stroke CT scans as normal and as stroke using a number of well-known pre-trained convolutional neural networks (CNNs), including GoogleNet, AlexNet, VGG-16, VGG-19, and Residual CNN. With an accuracy of 97.06%, VGG-19 obtained the highest score. Though the results are promising, they are limited to a specific and small dataset.
In conclusion, extensive research has been conducted on brain stroke detection using deep learning and CNN-based models. However, most existing studies rely on computationally expensive, heavyweight architectures and typically address only binary classification problems, differentiating stroke from non-stroke cases. This binary approach overlooks the clinical significance of distinguishing between different types of strokes, such as ischemic and hemorrhagic. To address these limitations, our study proposes a lightweight hybrid deep learning model designed to reduce computational complexity while maintaining high diagnostic accuracy. Additionally, we introduce a multi-class CT scan dataset comprising ischemic stroke, hemorrhagic stroke, and normal cases, which enables more clinically relevant and fine-grained classification.
@@ 3.3 Feature Extraction
In this research, we utilize pre-trained deep learning models, DenseNet201, InceptionV3, MobileNetV2, ResNet50, and Xception, to extract discriminative features from CT scan images. The process begins with image preprocessing, where images are resized to match the input dimensions required by each model (e.g., 224×224 pixels for DenseNet201, MobileNetV2, and ResNet50, and 299×299 pixels for InceptionV3 and Xception). The pixel values are then normalized between 0 and 1 to ensure numerical stability and faster convergence during model inference.
@@
3.3.1 MobileNetV2
3.3.2 DensNet201